using the method
subtitution method to solve simultaneous equations
x=y-2
x+y=18

Answer by tutorcecilia(2152)

x=y-2
x+y=18
.
(y-2)+y=18
2y-2=18
2y=18+2
2y=20
2y/2=20/2
y=10
.
x=y-2 [substitute y=10]
x=10-2
x=8
check by plugging in the values for x and y .